ROTATIONAL STUDIES IN THE HYDROXYBUTYRIC ACID SYSTEM

\underline{R. D. SUENRAM}, JASON PAJSKI, JUSTIN L. NEILL and BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., P.O. Box 400319, Charlottesville, VA 22904; MICHAEL J. TUBERGEN, Department of Chemistry, Kent State University, Kent, OH 44242.

The rotational spectra of α-, β-, and γ-hydroxybutyric acids have been investigated using CP-FTMW spectroscopy in order to study their conformational geometries and the effects of internal hydrogen bonding within the various compounds. Concurrently, ab initio calculations have been carried out in an effort to compare the theoretical and experimental results. In addition, the rotational spectrum of γ-butyrolactone has been observed so that similarities in the geometry can be compared to the conformational geometry of γ-hydroxybutyric acid. The γ-butyrolactone results when water is removed from γ-hydroxybutyric acid. This system is interesting because the two compounds can be interconverted. There is currently a good deal of interest in various methods of detecting and quantifying γ-hydroxybutyric acid in the forensic community. On the street, it is commonly known as GHB, a compound often used as a “date rape drug”.